# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

### v1

@@ 3.2 Consistency of subgroup estimation: Single rule setting
We now investigate the theoretical properties of CDT. We will show that under a set of regularity assumptions, CDT consistently recovers the optimal subgroups, and can improve the rate of convergence in estimating subgroups over standard decision trees. We show that the rate of convergence will depend on (1) the smoothness in the outcomes, and (2) how much noise there is in the underlying process. We focus our discussion on settings in which researchers are using standard greedy tree algorithms, such as CART, to estimate the second-stage decision tree. We note that in settings when researchers employ alternative algorithms to construct decision trees, some of the assumptions for consistency can be potentially relaxed.
Throughout, we denote $\mathbb{E}_{n}(A_{i})$ as the expectation of a variable $A_{i}$ over the observed sample (i.e., $\mathbb{E}_{n}(A_{i}):=\frac{1}{n}\sum_{i=1}^{n}A_{i}$), while $\mathbb{E}(A_{i})$ refers to the population-level expectation. Furthermore, for each unit $i$ and candidate split $k$, we define $\tau(X_{i},k)$ as the prediction generated from a decision tree, constructed with the true individual-level treatment effects $\tau_{i}$ and a single split at threshold $k$ (i.e., $\tau(X_{i},k):=\mathbbm{1}\{X_{i}\leq k\}\mathbb{E}(\tau_{i}\mid X_{i}\leq k)+\mathbbm{1}\{X_{i}>k\}\mathbb{E}(\tau_{i}\mid X_{i}>k)$). Similarly, we define $\tau^{d}(X_{i},k)$ as the prediction generated from a decision tree, constructed with the distilled individual-level treatment effects $\hat{\tau}_{i}^{d}$ and a single split at threshold $k$ (i.e., $\tau^{d}(X_{i},k):=\mathbbm{1}\{X_{i}\leq k\}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq k)+\mathbbm{1}\{X_{i}>k\}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}>k)$).
@@
An optimal split $s\in\mathcal{X}$ satisfies
$s:=\operatorname*{arg\,min}_{s^{\prime}\in\mathcal{X}}\mathbb{E}\left[\ell(\tau_{i};X_{i},s^{\prime})\right],$ (4)
where $\ell$ is defined as the squared loss function: $\ell(\tau_{i};X_{i},s^{\prime}):=(\tau_{i}-\tau(X_{i},s^{\prime}))^{2}.$
@@
Define $F_{X}$ to be the cumulative distribution function of the covariate $X$. Define the points $s_{l}$ and $s_{u}$ such that $10<F_{X}(s_{l})<F_{X}(s_{u})<$ 1. Let $s$ be an optimal split as defined in Definition 3.1.
  1. (a)
    Assume $F_{X}$ is absolutely continuous with a density $dF_{X}$ that is bounded away from zero and is continuous in a neighborhood of $s$.
  2. (b)
    Assume that $s$ is unique.
@@
Assumption (3.1) guarantees that there exists subgroups of interest to estimate. Intuitively, Assumption (3.1)-(a) ensures that the squared loss function will attain a minimum value for $s\in[s_{l},s_{u}]$. Assumption (3.1)-(b) rules out the possibility that there could be more than one potential split point. Furthermore, we are implicitly ruling out the setting considered in Cattaneo et al. (2022), in which the true optimal split does not exist.
@@
Given that the subgroups exist, we will now show that the estimated subgroups $\hat{\mathcal{G}}(X)$ from CDT converge to the true subgroups $\mathcal{G}(X)$. Greedy search guarantees that within a given sample of size $n$, the estimated split point, denoted $\hat{s}_{n}$, minimizes the within-sample loss (i.e., $\hat{s}_{n}:=\min_{s^{\prime}\in\text{supp}_{n}(X)}\mathbb{E}_{n}[\ell(\tau_{i};X,s^{\prime})]$). As such, in a hypothetical setting that researchers had access to $\tau_{i}$, the estimated subgroups $\hat{\mathcal{G}}(X)$ from a decision tree would give the optimal partitioning of the covariate space of $X$, within a given sample. Thus, to establish the consistency of CDT, we must consider two gaps. First, CDT uses the distilled treatment effects $\hat{\tau}_{i}^{d}$ in lieu of the individual-level treatment effect $\tau_{i}$. As such, we must constrain the differences that can arise from distilling the raw treatment effects with the teacher model. Second, we must show that the within-sample optimal partitioning will asymptotically recover the population-level optimal partitioning.
To address the first gap, we restrict our focus to valid teacher models, whose distilled treatment effects result in the same population-level optimal subgroups as the original treatment effects.
Assumption 3.2 (Valid Teacher Model).
Let $s$ denote the population-level optimal split according to (4), and define $s^{d}$ as
$s^{d}:=\operatorname*{arg\,min}_{s^{\prime}\in\mathcal{X}}\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X_{i},s^{\prime})\right].$ (5)
@@
Assume that the teacher model $\mathcal{M}$ used to distill the treatment effects into $\hat{\tau}^{d}_{i}$ satisfies $s^{d}=s$. In this case, we say that the teacher model is valid.
@@
Under validity of teacher model, the optimal partitioning of the covariate space of $X$ cannot change after distilling the individual-level treatment effects. In other words, the optimal population splits from the distilled individual-level treatment effects $\hat{\tau}_{i}^{d}$ will be equivalent to the optimal split with the true individual-level treatment effects $\tau_{i}$. A sufficient, though not necessary, condition for a valid teacher model is a correctly-specified teacher model, such that the residual noise that cannot be explained by the model must be uncorrelated with the covariates $X$ (i.e., exogeneity). In the context of a randomized control trial, we expect this to be true if the teacher model is flexible enough and the sample size is sufficiently large (details in Appendix A).
@@
We further assume that the amount that $\tau^{d}(X;s^{\prime})$ changes as the split point $s^{\prime}$ changes within a neighborhood of $s$ is bounded.
@@
Let $\tau_{0}=\frac{1}{2}\left\{\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}<s)+\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\geq s)\right\}$. Define $M(s)=\mathbb{E}\left[(\hat{\tau}^{d}_{i}-\tau_{0})1\{X_{i}\leq s\}\right]$. Let $\mathcal{N}$ be a neighborhood around $s$, where $\mathcal{N}:=(l,u)\subset[s_{l},s_{u}]$. Then, assume the following:
  • $\tau(x)$ is continuous in $(l,s)$ and $(s,u)$, where:
    $\lim_{x\to s^{-}}\tau(x)=\tau(s^{-})\text{ and }\lim_{x\to s^{+}}\tau(x)=\tau(s^{+})$
  • For $\epsilon>0$, there exists an $21<\alpha<$ 2 where:
    $\inf_{|s^{\prime}-s|<\epsilon,s\neq s^{\prime}}\frac{|M(s)-M(s^{\prime})|}{|s-s^{\prime}|^{\alpha}}>0$
  • There exists an $0<\eta\leq 1$ such that for all $\epsilon>0$:
    $\mathbb{E}\left[(\hat{\tau}^{d}_{i}-\tau_{0})^{2}\mathbbm{1}\{s-\epsilon\leq X_{i}\leq s+\epsilon\}\right]\leq C\epsilon^{2\eta}.$
Assumption (3.3) is a standard assumption in establishing consistency of $M$-estimators (e.g., Van der Vaart, 2000, Theorem 5.52). Informally, we can think of $\alpha$ and $\eta$ as parameters that control the smoothness of the underlying function around the optimal split point $s$. The assumption rules out settings in which $\hat{\tau}_{i}^{d}$ can take on values of infinity close to the optimal split point $s$. In practice, we expect Assumption (3.3) to generally hold across a large class of data generating processes. For example, in settings when $\tau^{d}(X;s^{\prime})$ is continuously differentiable in the neighborhood $\mathcal{N}$ and bounded in $s$, then $\alpha=2$ and $\eta=1/2$. Assumption (3.3) also allows for discontinuities at the split point $s$, so long as $\tau^{d}(X_{i};s)$ is continuous within the neighborhood around the split point.
Finally, we must assume bounded moments on the distilled treatment effects.
@@
Let $\hat{\mathcal{G}}_{1}(X_{i}):=\mathbbm{1}\{X_{i}\leq\hat{s}_{n}\}$ denote the subgroup estimated by CDT. Then, under Assumptions (3.1)-(3.4), CDT will consistently recover the subgroups:
@@
$\mathbb{E}\left[\left\lvert\hat{\mathcal{G}}_{1}(X_{i})-\mathcal{G}_{1}(X_{i})\right\rvert\right]\lesssim\left\lvert\hat{s}_{n}-s\right\rvert=O_{p}(n^{-1/2(\alpha-\eta)})$
@@
Proposition 3.1 formalizes that the rate of convergence depends directly on how smooth the underlying data generating process is. The rate of convergence in the estimated subgroups inherits a characteristic highlighted in Escanciano (2020), which bridged the gap in convergence rate of the estimated split in a decision tree $\hat{s}_{n}$ between a fully continuous setting (considered in Banerjee and McKeague, 2007; Buhlmann and Yu, 2002), and discontinuous settings, as considered in Chan (1993) and Kosorok (2008).
@@
Consider the setting in which $\tau^{d}(X;s^{\prime})$ is continuously differentiable. This would imply a rate of convergence of $O_{p}(n^{-2/3})$. On the other hand, in the setting when $\tau(X)$ is discontinuous at $X=s$, then $\hat{s}-s=O_{p}(n^{-1})$. Intuitively, as the underlying data generating process becomes more discontinuous at the split point $s$, the faster the estimated split $\hat{s}_{n}$ will converge to the true split $s$. While these results have been studied in the decision tree and change point estimation literature (see, for example, Banerjee and McKeague, 2007; Buhlmann and Yu, 2002; Escanciano, 2020), we are, to the best of our knowledge, the first to formally connect the notion of split consistency to subgroup recovery.
@@
Proposition 3.1 provides practical guidance in considering what suitable teacher models should be used to best improve the rate of convergence using distillation. In standard decision tree settings, $\alpha$ and $\eta$ are something inherent to the underlying data generating process. In contrast, under distillation, we are considering the underlying smoothness of not the original outcomes $\tau_{i}$, but the distilled $\hat{\tau}_{i}^{d}$, which are constructed from a teacher model. This implies that in settings when the underlying $\tau(X)$ is smooth around the optimal split point, but the teacher model correctly constructs piecewise functions around the partitioning points, then this can improve the rate of convergence from $O_{p}(n^{-2/3})$ to the much faster rate of $O_{p}(n^{-1})$.
@@
In addition to consistently recovering the subgroups, CDT can further improve the stability of the split estimation by de-noising the individual-level treatment effect and increasing the signal strength in the individual-level treatment effect. To illustrate, in the following example we consider a setting in which researchers have access to the true individual-level treatment effect $\tau_{i}$ and estimate a decision tree using the covariates $X_{i}$. In practice, this is infeasible, as researchers can never observe $\tau_{i}$, but this comparison serves as a helpful benchmark to compare the potential improvements from distillation.
@@
Consider a setting where $\tau(X)$ is continuously differentiable and bounded in a neighborhood around $s$. Furthermore, assume we can write the individual-level treatment effect and the distilled treatment effects as a function of $\tau(X_{i})$ and a noise term:
$\tau_{i}=\tau(X_{i})+v_{i},\text{ and }\hat{\tau}_{i}^{d}=\tau(X_{i})+v_{i}^{d}.$
Assume that $\mathbb{E}(v_{i})=\mathbb{E}(v_{i}^{d})=0$, both $\text{var}(v_{i})$ and $\text{var}(v_{i}^{d})$ are finite, and $\text{var}(\tau(X_{i}))>0$. Then, under squared loss and mild regularity assumptions (i.e., Assumption (B.3)), the relative asymptotic variance of the estimated splits without distillation $\hat{s}_{n}$ and with distillation $\hat{s}_{n}^{d}$ is a function of the relative signal-to-noise ratio between the individual-level treatment effect and the distilled individual-level treatment effects:
$\frac{\text{asyvar}(\hat{s}_{n})}{\text{asyvar}(\hat{s}_{n}^{d})}=\left(\frac{\text{SNR}_{distil}}{\text{SNR}_{original}}\right)^{2+4/3},$
@@
The results from Example 3.1 follow from noting that the asymptotic distribution of the estimated splits in a linear setting follow Chernoff’s distribution (Buhlmann and Yu, 2002; Groeneboom, 1989). While the variance itself is not straightforward to interpret and depends on an Airy function, the ratio of variances simplifies into a function of the signal-to-noise ratios. See Appendix B for the full derivation.
Example 3.1 formalizes the intuition that as the first-stage learner more closely approximates $\tau(X)$ (i.e., the conditional expectation function of $\tau$, given $X$), the stability of the estimated splits under distillation also improves. In particular, we can view the first-stage learner as a de-noising, or smoothing, step that filters the original individual-level treatment effect $\tau_{i}$ to a less noisy representation, in the space of $X$.
3.3 Consistency of subgroup estimation: General setting with multiple rules
We next extend the consistency results from the previous subsection to the general multivariate setting, in which we observe $p$ total covariates (i.e., $X\in\mathbbm{R}^{p}$) and the optimal subgroups, satisfying (3), consist of an arbitrary number of rules with different covariates and different split points. Throughout, we will denote the total number of binary decision rules for a subgroup as $r_{g}$. In this setting, we can rewrite the optimal subgroups in terms of their binary decision rules. For example, without loss of generality, the first subgroup $\mathcal{G}_{1}(X)$ would be written as $\mathcal{G}_{1}(X)=\prod_{k=1}^{r_{1}}\mathbbm{1}\left\{X^{(j_{k})}\leq s^{(j_{k})}\right\},$ where $s^{(j_{k})}\in\text{supp}(X^{(j_{k})})$, and $j_{1},\ldots,j_{r_{1}}\in\{1,\ldots,p\}$. Under this setting, estimating consistent subgroups requires not only consistently recovering the correct split points, but also selecting the right covariates to split on.
To show this consistency, we must introduce an additional separability condition, which states that conditional on the previously selected binary decision rules, the differences in loss from constructing a relevant and irrelevant rule must be greater than zero (i.e., separable).

### v2

@@ 3.2 Consistency of subgroup estimation: Single rule setting
We now investigate the theoretical properties of CDT. We show that under a set of regularity assumptions, CDT consistently recovers the optimal subgroups, and can improve the rate of convergence in estimating subgroups over standard decision trees. The rate of convergence depends on (1) the smoothness in the outcomes, and (2) how much noise there is in the underlying process. We focus our discussion standard greedy tree algorithms, such as CART. In settings when researchers employ alternative algorithms to construct decision trees, some of the assumptions for consistency can be potentially relaxed.
Throughout, we denote $\mathbb{E}_{n}(A_{i})$ as the expectation of a variable $A_{i}$ over the observed sample (i.e., $\mathbb{E}_{n}(A_{i}):=\frac{1}{n}\sum_{i=1}^{n}A_{i}$), while $\mathbb{E}(A_{i})$ refers to the population-level expectation. Furthermore, for each unit $i$, we define $f_{\tau}(X_{i};k)$ as the prediction generated from a decision tree estimated with $\tau_{i}$, at a single split with threshold $k$ (i.e., $f_{\tau}(X_{i};k):=\mathbbm{1}\{X_{i}\leq k\}\mathbb{E}(\tau_{i}\mid X_{i}\leq k)+\mathbbm{1}\{X_{i}>k\}\mathbb{E}(\tau_{i}\mid X_{i}>k)$). Similarly, we define $f^{d}_{\tau}(X_{i};k)$ as the prediction generated from a decision tree using $\hat{\tau}_{i}^{d}$, at a a single split with threshold $k$ (i.e., $f^{d}_{\tau}(X_{i};k):=\mathbbm{1}\{X_{i}\leq k\}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\leq k)+\mathbbm{1}\{X_{i}>k\}\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}>k)$).
@@
An optimal split $s\in\mathcal{X}$ satisfies $s:=\operatorname*{arg\,min}_{s^{\prime}\in\mathcal{X}}\mathbb{E}\left[\ell(\tau_{i};X_{i},s^{\prime})\right],$ where $\ell$ is defined as the squared loss function: $\ell(\tau_{i};X_{i},s^{\prime}):=(\tau_{i}-f_{\tau}(X;s^{\prime}))^{2}.$
@@
Let $F_{X}$ be the cumulative distribution function of the covariate $X$. Define the points $s_{l}$ and $s_{u}$ such that $10<F_{X}(s_{l})<F_{X}(s_{u})<$ 1. Let $s$ be an optimal split as defined in Definition 3.1. Assume that (i) $F_{X}$ is absolutely continuous with a density $dF_{X}$ that is bounded away from zero and is continuous in a neighborhood of $s$, and (ii) $s$ is unique.
@@
Assumption (3.1) guarantees that there exists subgroups of interest to estimate. Intuitively, the constraint on $F_{X}$ ensures that the squared loss function will attain a minimum value for $s\in[s_{l},s_{u}]$. Furthermore, Assumption (3.1) rules out the possibility that there could be more than one potential split point. In particular, we are implicitly ruling out the setting considered in Cattaneo et al. (2022), in which an optimal split does not exist.
@@
Given that the subgroups exist, we next show that the estimated subgroups $\hat{\mathcal{G}}(X)$ from CDT converge to the true subgroups $\mathcal{G}(X)$. To do so, recall that greedy decision tree algorithms guarantee that within a given sample of size $n$, the estimated split point, denoted $\hat{s}_{n}$, minimizes the within-sample loss (i.e., $\hat{s}_{n}:=\min_{s^{\prime}\in\text{supp}_{n}(X)}\mathbb{E}_{n}[\ell(\tau_{i};X,s^{\prime})]$). Thus, in the hypothetical setting where researchers had access to $\tau_{i}$, the estimated subgroups $\hat{\mathcal{G}}(X)$ from a decision tree would give the optimal partitioning of the covariate space of $X$, within a given sample. To hence establish the consistency of CDT, we must consider two gaps. First, CDT uses the predicted treatment effects $\hat{\tau}_{i}^{d}$ in lieu of the individual-level treatment effect $\tau_{i}$. As such, we must constrain the differences that can arise from distilling the raw treatment effects with the teacher model. Second, we must show that the within-sample optimal partitioning will asymptotically recover the population-level optimal partitioning.
To address the first gap, we restrict our focus to valid teacher models.
Assumption 3.2 (Valid Teacher).
@@
Let $s$ be the population-level optimal split, and define $s^{d}:=\operatorname*{arg\,min}_{s^{\prime}\in\mathcal{X}}\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X_{i},s^{\prime})\right]$. Assume the teacher model $\mathcal{M}$ used to distill the treatment effects into $\hat{\tau}^{d}_{i}$ satisfies $s^{d}=s$. In this case, we say that the teacher model is valid.
@@
Under this validity of the teacher model, the population-level optimal subgroups remain unchanged after distilling the individual-level treatment effects. In other words, the optimal population splits from the predicted individual-level treatment effects $\hat{\tau}_{i}^{d}$ are equivalent to the optimal split with the true individual-level treatment effects $\tau_{i}$. A sufficient, though not necessary, condition for a valid teacher model is a correctly-specified teacher model, such that the residual noise that cannot be explained by the model must be uncorrelated with the covariates $X$ (i.e., exogeneity). In randomized control trials, we expect this to be true if the teacher model is flexible enough and the sample size is sufficiently large. Assumption (3.2) can also be relaxed to consider approximately valid teacher models (see Appendix A.2 for additional discussion).
@@
We further assume that the amount that $f^{d}_{\tau}(X;s^{\prime})$ changes as the split point $s^{\prime}$ changes within a neighborhood of $s$ is bounded.
@@
Let $\tau_{0}=\frac{1}{2}\left\{\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}<s)+\mathbb{E}(\hat{\tau}^{d}_{i}\mid X_{i}\geq s)\right\}$. Define $H(s)=\mathbb{E}\left[(\hat{\tau}^{d}_{i}-\tau_{0})1\{X_{i}\leq s\}\right]$. Let $\mathcal{N}$ be a neighborhood around $s$, where $\mathcal{N}:=(l,u)\subset[s_{l},s_{u}]$. Then, assume the following:
  • $\tau(x)$ is continuous in $(l,s)$ and $(s,u)$, where $\lim_{x\to s^{-}}\tau(x)=\tau(s^{-})\text{ and }\lim_{x\to s^{+}}\tau(x)=\tau(s^{+})$.
  • For $\epsilon>0$, there exists an $21<\alpha<$ 2 where: $\inf_{|s^{\prime}-s|<\epsilon,s\neq s^{\prime}}\frac{|H(s)-H(s^{\prime})|}{|s-s^{\prime}|^{\alpha}}>0$.
  • There exists $0<\eta\leq 1$ such that $\forall~{}\epsilon>0$, $\mathbb{E}\left[(\hat{\tau}^{d}_{i}-\tau_{0})^{2}\mathbbm{1}\{s-\epsilon\leq X_{i}\leq s+\epsilon\}\right]\leq C\epsilon^{2\eta}.$
Assumption (3.3) is a standard assumption in establishing consistency of $M$-estimators (e.g., Van der Vaart, 2000, Theorem 5.52). Informally, $\alpha$ and $\eta$ are parameters that control the smoothness of the underlying function around the optimal split point $s$. The assumption rules out settings where $\hat{\tau}_{i}^{d}$ can take on values of infinity close to the optimal split point $s$. In practice, we expect Assumption (3.3) to generally hold across a large class of data generating processes. For example, when $\hat{\tau}_{i}^{d}$ is continuously differentiable in the neighborhood $\mathcal{N}$ and bounded in $s$, then $\alpha=2$ and $\eta=1/2$. Assumption (3.3) also allows for discontinuities at the split point $s$, so long as $\hat{\tau}_{i}^{d}$ is continuous within the neighborhood around the split point.
Finally, we assume the predicted treatment effects have bounded moments.
@@
Let $\hat{\mathcal{G}}_{1}(X_{i}):=\mathbbm{1}\{X_{i}\leq\hat{s}_{n}\}$ denote the subgroup estimated by CDT. Then, under Assumptions (3.1)-(3.4), CDT will consistently recover the optimal subgroups:
@@
$\mathbb{E}\left[\left\lvert\hat{\mathcal{G}}_{1}(X_{i})-\mathcal{G}_{1}(X_{i})\right\rvert\right]\lesssim\left\lvert\hat{s}_{n}-s\right\rvert=O_{p}(n^{-1/2(\alpha-\eta)}).$
@@
Proposition 3.1 formalizes that the rate of convergence depends directly on how smooth the underlying data generating process is. The rate of convergence in the estimated subgroups inherits a characteristic highlighted in Escanciano (2020), which bridged the gap in convergence rate of the estimated split in a decision tree $\hat{s}_{n}$ between a fully continuous setting (considered in (Banerjee & McKeague, 2007; Buhlmann & Yu, 2002)) and discontinuous settings (considered in (Chan, 1993; Kosorok, 2008)).
@@
Consider the setting in which $\hat{\tau}_{i}^{d}$ is continuously differentiable. This would imply a convergence rate of $O_{p}(n^{-2/3})$. On the other hand, when $\hat{\tau}_{i}^{d}$ is discontinuous at $X=s$, then $\hat{s}-s=O_{p}(n^{-1})$. Intuitively, as the underlying data generating process becomes more discontinuous at the split point $s$, the faster the estimated split $\hat{s}_{n}$ will converge to the true split $s$. While these results have been studied in the decision tree and change point estimation literature (see, for example, Banerjee & McKeague, 2007; Buhlmann & Yu, 2002; Escanciano, 2020), we are, to the best of our knowledge, the first to formally connect the notion of split consistency to subgroup recovery.
@@
Proposition 3.1 provides practical guidance in considering what suitable teacher models should be used to best improve the rate of convergence using distillation. In standard decision tree settings, $\alpha$ and $\eta$ are something inherent to the underlying data generating process. In contrast, under distillation, we are considering the underlying smoothness of not the original outcomes $\tau_{i}$, but the distilled $\hat{\tau}_{i}^{d}$, which are constructed from a teacher model. This implies that when the underlying $\tau(X)$ is smooth around the optimal split point but the teacher model correctly constructs piecewise functions around the partitioning points, then the convergence rate can improve from $O_{p}(n^{-2/3})$ to the much faster rate of $O_{p}(n^{-1})$.
@@
In addition to consistently recovering the subgroups, CDT can further improve the stability of the split estimation by de-noising the individual-level treatment effect and increasing the signal strength in the individual-level treatment effect. To illustrate, in the following example, we consider a setting in which researchers have access to the true individual-level treatment effect $\tau_{i}$ and estimate a decision tree using the covariates $X_{i}$. In practice, this is infeasible, as researchers can never observe $\tau_{i}$, but this comparison serves as a helpful benchmark to compare the potential improvements from distillation.
@@
Consider a setting where $\tau(X)$ is continuously differentiable and bounded in a neighborhood around $s$. Furthermore, assume we can write the individual-level treatment effect and the predicted treatment effects as a function of $\tau(X_{i})$ and a noise term (i.e., $\tau_{i}=\tau(X_{i})+v_{i}$, and $\hat{\tau}_{i}^{d}=\tau(X_{i})+v_{i}^{d}$), where $\mathbb{E}(v_{i})=\mathbb{E}(v_{i}^{d})=0$, both $\text{var}(v_{i})$ and $\text{var}(v_{i}^{d})$ are finite, and $\text{var}(\tau(X_{i}))>0$. Then, under squared loss and mild regularity assumptions (Assumption (B.3)), the relative asymptotic variance of the estimated splits without distillation $\hat{s}^{orig}_{n}$ and with distillation $\hat{s}_{n}$ is
$\frac{\text{asyvar}(\hat{s}^{orig}_{n})}{\text{asyvar}(\hat{s}_{n})}=\left(\frac{\text{SNR}_{distil}}{\text{SNR}_{original}}\right)^{2+4/3},$
@@
The results from Example 3.1 follow from noting that the asymptotic distribution of the estimated splits in a linear setting follow Chernoff’s distribution (Buhlmann & Yu, 2002; Groeneboom, 1989). While the variance itself is not straightforward to interpret and depends on an Airy function, the ratio of variances simplifies into a function of the signal-to-noise ratios (derivation in Appendix B). Example 3.1 formalizes the intuition that as the first-stage learner more closely approximates $\tau(X)$ (i.e., the conditional expectation function of $\tau_{i}$, given the pre-treatment covariates), the stability of the estimated splits after distillation also improves. We can hence view the first-stage learner as a de-noising, or smoothing, step that filters the original individual-level treatment effect $\tau_{i}$ to a less noisy representation, in the space of $X$.
3.3 General setting with multiple rules
We next extend the consistency results to the multivariate setting, in which we observe $p$ total covariates (i.e., $X\in\mathbbm{R}^{p}$) and the optimal subgroups, satisfying (3), consist of an arbitrary number of rules with different covariates and different split points. Throughout, for each subgroup $\mathcal{G}_{g}(X)$, we denote the total number of binary decision rules as $r_{g}$.
Under this multivariate setting, estimating consistent subgroups requires not only recovering the correct split points, but also selecting the right covariates to split on. We thus must introduce an additional separability condition, which states that conditional on the previously selected binary decision rules, the differences in loss from constructing a relevant and irrelevant rule must be greater than zero (i.e., separable).